OpenStudy (anonymous):

The perimeter of some traffic must be 110 inches its length is 13 inches longer than its width. Find the dimensions of this sign The width of the sign is __inch The length of the sign is ____ inch

OpenStudy (anonymous):

2L +2W = P L=W+13 2(W+13) + 2W=P P=110

OpenStudy (anonymous):

width and length are what i don't understand